OpenStudy (crashonce):

another GP problem @ganeshie8

OpenStudy (crashonce):

\[1+3^{2}+3^{4}+...+ 3^{2n}\]

ganeshie8 (ganeshie8):

how many terms are there in the series ?

OpenStudy (crashonce):

it just asks you to add up to \[3^{2n}\]

ganeshie8 (ganeshie8):

lets count the number of terms again

ganeshie8 (ganeshie8):

Notice the exponents : 0, 2, 4, ... , 2n

ganeshie8 (ganeshie8):

how many terms are they ?

OpenStudy (crashonce):

2n+1

OpenStudy (crashonce):

@ganeshie8

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

0, 2, 4, ... , 2n if i divide each term by 2, the number of terms wont change, right ?

ganeshie8 (ganeshie8):

so lets divide by 2, you get : 0, 1, 2, ... , n

ganeshie8 (ganeshie8):

how many terms are they ?

OpenStudy (crashonce):

n+1

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

and r = 3^2 as you said earlier

ganeshie8 (ganeshie8):

use the partial sum formula : \[\rm S_k = a_1 \dfrac{r^k-1}{r-1}\]

ganeshie8 (ganeshie8):

r = 3^2 k = n+1

OpenStudy (crashonce):

ok i got it thanks!!!

OpenStudy (crashonce):

ive got one more sorry